# SLoRD: Structural Low-Rank Descriptors for Shape Consistency in Vertebrae Segmentation

### v1

@@ Abstract
Automatic and precise segmentation of vertebrae from CT images is crucial for various clinical applications. However, due to a lack of explicit and strict constraints, existing methods especially for single-stage methods, still suffer from the challenge of intra-vertebrae segmentation inconsistency, which refers to multiple label predictions inside a singular vertebra. For multi-stage methods, vertebrae detection serving as the first step, is affected by the pathology and mental implants. Thus, incorrect detections cause biased patches before segmentation, then lead to inconsistent labeling and segmentation. In our work, motivated by the perspective of instance segmentation, we try to label individual and complete binary masks to address this limitation. Specifically, a contour-based network is proposed based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. These contour descriptors are acquired in a data-driven manner in advance. For a more precise representation of contour descriptors, we adopt the spherical coordinate system and devise the spherical centroid. Besides, the contour loss is designed to impose explicit consistency constraints, facilitating regressed contour points close to vertebral boundaries. Quantitative and qualitative evaluations on VerSe 2019 demonstrate the superior performance of our framework over other single-stage and multi-stage state-of-the-art (SOTA) methods. Source codes are available https://github.com/AlexYouXin/SLoRD-VerSe.
@@ 1 Introduction
Automatic segmentation of vertebrae from CT images plays a significant role in spinal disease treatment, including preoperative diagnosis, surgical guidance, and postoperative assessment [1, 2, 3]. Moreover, the shape analysis of vertebrae serves as an important anatomical reference for other anatomies in clinical practices [4, 5]. According to the anatomical prior of vertebral structures, it is emerging to preserve the consistency of label predictions inside a singular vertebra [6]. However, as illustrated in Figure 1, for single-stage and multi-stage methods, predicted masks for the specific vertebra might contain inconsistent label predictions, and partial voxel predictions in adjacent vertebrae reveal the same label. That phenomenon has been recognized as intra-vertebrae label inconsistency [7], which poses a large challenge for the task of vertebrae segmentation.
@@
Specifically, earlier methods solved vertebrae segmentation from CT images by employing statistical models [8], active shape models [9], level sets [10] and graph cuts theories [11]. However, those approaches fail to generalize well on unseen data. Recent deep learning-based methods demonstrate a more promising performance for vertebrae image segmentation even with arbitrary FOVs. One category of existing works [1, 12, 7] adopted single-stage models. Lessmann et al. [1] used fully convolutional networks by iteratively analyzing image patches for automatic vertebrae segmentation and identification. However, this method relies on the perfect prediction for the first label and hardly recovers from failures in the top vertebra. You et al. [7] proposed an edge detection block to restrain boundaries for better consistency inside the boundary box. However, these models still suffer from drawbacks of similar appearances between adjacent vertebrae and the patch-based inference paradigm, resulting in intra-vertebrae label inconsistency. The other methods [13, 14, 15, 16] designed a multi-stage pipeline to improve the segmentation performance. Tao et al. [15] introduced a Transformer to handle the labeling task, then employed 3D UNet [17] with the original CT image and an auxiliary heatmap as inputs to segment all vertebrae. Besides, Payer et al. [16] implemented a three-step approach to first localize the spine, then simultaneously located and identified each vertebra, and finally segmented each vertebra. However, a biased detected patch due to the inaccurate center regression will enforce inconsistency inside vertebrae, particularly for pathological vertebrae with abnormal shapes or mental implants as illustrated by Figure 1(d & e). Also, this limitation can be referred to according to the qualitative analysis of experimental results in [15].
Carrying out the cause and effect analysis, we summarize three leading factors contributing to the intra-vertebrae segmentation inconsistency for CT images. ❶ Neighbored vertebrae show similar appearances and intensity values from CT images, which results in label confusion between adjacent vertebrae [18]. ❷ Vertebrae CT scans display arbitrary field-of-views (FOVs) with various label distributions [19]. Consequently, U-shape networks [20, 21] suffer from a poor identification of the spine, especially for bi-terminal vertebrae [7]. ❸ A portion of spines bear pathological vertebrae with abnormal or deformed shapes [13]. And there also exist metal artifacts in some CT scans. In brief, vertebral data show different characteristics from other anatomical structures, making it difficult to precisely and consistently segment vertebrae with sequential labels.
@@
Figure 1: (a) CT case 1. (b-c) Coarse prediction by MedNeXt [22] & refined outcome by MedNeXt + SLoRD on case 1. (d) Case 2 with meta artifacts. (e-f) Coarse prediction by the multi-stage pipeline Spine-Transformer [15] & our refined prediction on case 2. (g) Spherical coordinate system for 3D vertebral contour descriptions. x, y and z axis represent voxel orientations from Right/Anterior/Inferior to Left/Posterior/Superior. $\rho$: the radial distance away from the spherical center, $\phi$: the polar angle between the radial vector and $z^{+}$, $\theta$: the azimuth angle between $x^{+}$ and the projected radial vector in the XOY plane.
@@
Essentially, all the methods above do not impose explicit and strict constraints for voxels inside each vertebra during training, resulting in label inconsistency. To address this challenge, motivated by the concept of instance segmentation [1], we attempt to label individual and complete binary masks to achieve multi-label vertebrae segmentation. Due to the connection between superior and inferior articular processes [2], the main challenge transforms to how to divide joint binary vertebral masks without destroying original contours. Thus, we devise an effective contour-based network based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. Specifically, vertebrae show similar 3D shapes according to anatomical structures, lumbar and thoracic vertebrae in particular. Hence, we can find out a low-dimension contour space to represent all vertebrae data. Intuited by [23], we build a novel representation way of contours based on the spherical coordinate system as shown in Figure 1(g). Additionally, spherical centroids are proposed to enhance the precision of contour representation vectors. Via a linear combination between basic vectors of the contour space, instance masks of vertebrae can be approximately restored. Then vertebral labels provided by coarse predictions facilitate the generation of semantic masks. Finally, we incorporate this pipeline into deep neural networks to boost the framework’s generalization abilities on unseen datasets. And a contour loss is devised to exert a consistency constraint for each vertebra. Extensive experiments are carried out on the public VerSe 2019 dataset [18]. And our method achieves SOTA performance compared with other single-stage and multi-stage approaches. Moreover, qualitative results demonstrate that our two-stage pipeline can significantly improve the intra-vertebrae segmentation inconsistency.
@@ 2 Methodology
2.1 Preliminaries on Contour Descriptors
For designing contour descriptors, we incorporate the idea from PolarMask [23, 24], in which each contour point is encoded with the polar coordinate. Specifically, the origin point is set as the centroid of the object, and each contour point is determined by the fixed sampling angle and corresponding distance. Directional angles from $0^{\circ}$ to $360^{\circ}$ efficiently connect each point into a complete contour. To strike a balance between the precision and computational cost, the polar system makes good use of the non-convex shape property of vertebrae anatomies and chooses the maximum distance at the same angle.
Since it is challenging to extend the polar system to contour representations of 3D objects like vertebrae, we formulate contours with the spherical coordinate system as shown in Figure 1(g). With respect to the inner centroid, we describe the boundary using spherical coordinates $(\rho_{{}_{\mathrm{ij}}},\theta_{i},\phi_{j}),i=1,...,I;j=1,...,J$. Here $\rho$ represents the length of the radial vector from the spherical center to the contour point, $\phi$ is the polar angle between the radial vector and $z^{+}$, $\theta$ means the azimuth angle between $x^{+}$ and the projected radial vector in the XOY plane. Angular coordinates $\theta_{i},\phi_{j}$ are sampled uniformly with a sampling interval of $5^{\circ}$, with I equal to $\tfrac{360^{\circ}}{5^{\circ}}=72$, J equal to $(\tfrac{180^{\circ}}{5^{\circ}}+1)=37$. Thus, given the sampling angle grid, only radial coordinates are recorded to represent the contour:
@@ 2.1 Preliminaries on Contour Descriptors
Here $\bm{\rho}$ refers to the contour descriptor, with the dimension $N$ equal to $I\times J=2664$. Since shapes of most vertebrae are well structured and resembling to each other, low-rank contour descriptors can be exploited in a data-driven manner via Singular Value Decomposition (SVD) [24]. Detailedly, a non-convex contour matrix $\mathcal{M}=[\bm{\rho}_{1},\cdot\cdot\cdot,\bm{\rho}_{L}]$ is constructed from L vertebrae in the training data. Then we perform SVD on matrix $\mathcal{M}$:
@@
Where $\bm{U}^{N\times N}=[\bm{u}_{1},\cdot\cdot\cdot,\bm{u}_{N}]$ and $\bm{V}^{L\times L}=[\bm{v}_{1},\cdot\cdot\cdot,\bm{v}_{L}]$ are orthogonal matrixes and $\Sigma^{N\times L}$ is a rectangular diagonal matrix, consisting of singular values $[\sigma_{1},\cdot\cdot\cdot,\sigma_{r}]\,(r=\min(N,L)$. The best rank-k approximation of $\mathcal{M}$ can be calculated as follows:
$\mathcal{M}(k)=\sigma_{1}\bm{u}_{1}\bm{v}_{1}^{T}+\cdot\cdot\cdot+\sigma_{k}\bm{u}_{k}\bm{v}_{k}^{T}=[\bm{u}_{1},\cdot\cdot\cdot,\bm{u}_{k}]\,\mathcal{C}=\bm{U}(k)\,\mathcal{C}\approx\mathcal{M}$ (3)
$\mathcal{C}=[\sigma_{1}\bm{v}_{1},\cdot\cdot\cdot,\sigma_{k}\bm{v}_{k}]^{T}=\bm{U}(k)^{T}\mathcal{M}$ (4)
Here $\bm{U}(k)^{N\times k}$ refers to k basic contour descriptors, which can span into the contour space and be acquired from the training data in advance. $\mathcal{C}^{k\times L}$ means combinational coefficients, calculated by eigen descriptors and matrix $\mathcal{M}$. As long as $\mathcal{C}$ is correct, we can restore the original contours of vertebrae approximately as illustrated by Figure 2. For a better trade-off between performance and efficiency, only the first 200 basic contour descriptors are adopted in our work.
Figure 3: The two-stage framework. (a) Coarse predictions by arbitrary segmentation networks. (b) Based on the combination of basic contour descriptors, SLoRD aims to regress contour points close to vertebral boundaries. (c) Iterative refinement for coarse masks.
@@ 2.2 Spherical Centroid
Although the encoding scheme for vertebral contours is efficient, it is very tricky to determine the spherical center, which will significantly affect the restoration performance of vertebral shapes. Since the polar angle $\phi$ and azimuth angle $\theta$ are both sampled at equal intervals, a spherical center with a biased location will result in fewer sampling contour points for the boundary region far from it. Consequently, that region suffers from a contour under-representation. Hence, it has been a crucial topic on how to designate the spherical center.
@@
$C_{s}=\mathop{\arg\min}\limits_{c\in M}(\tfrac{1}{|\mathcal{B}|}\displaystyle\sum_{b\in\mathcal{B}}d\,(c,b)+\lambda\cdot d\,(c,\delta)),\ \mathrm{s.t.}\ \delta=\mathop{\arg\max}\limits_{j\in\mathcal{B}}j_{y}$ (6)
@@
Where $\mathcal{F}(\cdot)$ represents the morphological dilation operation, $M$ refers to the segmentation mask of each vertebral volume, the boundary map $\mathcal{B}$ is the difference set between the dilated mask $\mathcal{F}(M)$ and the mask itself $M$. Besides, $d(\cdot)$ means the Euclidean distance. The first item in Eq (6) refers to the average distance from each boundary point $b$ to the center $c\in\mathcal{P}$. Besides, considering the non-convex shape prior of the spinous process, we attempt to push the spherical centroid closer to it. That will enlarge the value range of $\theta$ and $\phi$, then enhance the shape representations of the spinous process. That is why the second regularization term is incorporated, in which $\delta$ is the boundary point with the largest y-dimensional index. $\lambda$ is a scaling parameter, empirically set as $0.005$.
2.3 Contour Generation with Contour Priors

### v2

@@ Abstract
Automatic and precise multi-class vertebrae segmentation from CT images is crucial for various clinical applications. However, due to a lack of explicit consistency constraints, existing methods especially for single-stage methods, still suffer from the challenge of intra-vertebrae segmentation inconsistency, which refers to multiple label predictions inside a singular vertebra. For multi-stage methods, vertebrae detection serving as the first step, tends to be affected by the pathology and metal implants. Thus, imprecise detections cause biased patches before segmentation, which then leads to inaccurate contour delineation and inconsistent segmentation. In our work, we intend to label individual and complete binary masks to address that challenge. Specifically, a contour generation network is proposed based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For a structural representation of vertebral contours, we adopt the spherical coordinate system and devise the spherical centroid to calculate contour descriptors. Due to vertebrae’s similar appearances, basic contour descriptors can be acquired to restore original contours. Therefore, SLoRD leverages these contour priors and explicit shape constraints to facilitate regressed contour points close to vertebral surfaces. Quantitative and qualitative evaluations on VerSe 2019 and 2020 demonstrate the superior performance of our framework over other single-stage and multi-stage state-of-the-art (SOTA) methods. Further, SLoRD is a plug-and-play framework to refine the segmentation inconsistency existing in coarse predictions from other approaches. Source codes are available https://github.com/AlexYouXin/SLoRD-VerSe.
{IEEEkeywords}
Vertebrae segmentation, Inconsistency, Contour descriptors, Refinement
@@ 1 Introduction
Automatic multi-class vertebrae segmentation from CT images plays a significant role in spinal disease treatment, including preoperative diagnosis, surgical guidance, and postoperative assessment [22, 5, 21]. Moreover, the shape analysis of vertebrae serve as an important anatomical reference for other anatomies in clinical practices [9, 28]. According to the anatomical prior of vertebral structures, it is emerging to preserve the consistency of label predictions inside a singular vertebra [24]. However, as illustrated in Fig.1, for single-stage and multi-stage methods, predicted masks for the specific vertebra might contain inconsistent label predictions, and partial voxel predictions in adjacent vertebrae reveal the same label. That phenomenon has been recognized as intra-vertebrae label inconsistency [45], which poses an enormous challenge for the task of precise vertebrae segmentation.
Figure 1: (a) CT case 1. (b-c) Coarse prediction by 3D UNet & refined outcome by 3D UNet + SLoRD on case 1. (d) CT case 2 with metal artifacts. (e-f) Coarse prediction by the multi-stage pipeline Spine-Transformer [38] & our refined prediction on case 2.
@@
Specifically, earlier methods solved vertebrae segmentation from volumetric CT images by employing statistical models [20], active shape models [14], level sets [23], Markov random field [17, 18], graph cuts [3] and machine learning theories [10, 19]. However, those approaches fail to perform well on unseen data due to insufficient generalization ability. Current methods based on deep learning techniques demonstrate a more promising performance for vertebrae image segmentation even with arbitrary field-of-views (FOVs). One category of existing works [11, 15, 45] adopted single-stage convolutional or Transformer-based networks for automatic vertebrae segmentation and identification. However, these models suffer from drawbacks of similar appearances between adjacent vertebrae and the patch-based inference paradigm [22], resulting in intra-vertebrae label inconsistency as revealed in Fig.1(a & b). The other methods [27, 26, 38, 31] designed a multi-stage pipeline, which can boost the segmentation performance by reducing the disturbance from neighboring contexts based on the localize-then-segment concept. However, a biased local patch from the detection network due to the inaccurate center regression will enforce inconsistency inside vertebrae, particularly for pathological vertebrae with abnormal shapes or metal implants as illustrated by Fig.1(d & e). Also, this limitation can be referred to according to the qualitative analysis of experimental results in [38].
Essentially, all the above methods fail to explicitly impose both contour precision and intra-vertebrae voxel consistency constraints synchronously, resulting in segmentation inconsistency. To resolve this limitation, we attempt to label individual and complete binary masks to achieve multi-label vertebrae segmentation. Due to the connection between superior and inferior articular processes [5], the main challenge is transformed to how to divide joint binary vertebral masks while preserving original contours to the greatest extent [13].
@@
Figure 2: Spherical coordinate system for 3D vertebral contour descriptions. x, y, and z axis represent voxel orientations from Right/Anterior/Inferior to Left/Posterior/Superior. $\rho$: the radial distance away from the spherical center, $\phi$: the azimuth angle between the radial vector and $z^{+}$, $\theta$: the polar angle between $x^{+}$ and the projected radial vector in the XOY plane.
@@
Thus, we devise an effective contour generation network based on Structural Low-Rank Descriptors for shape consistency, termed SLoRD. For the contour representation, the complete point cloud of vertebral boundaries shows an unorganized arrangement with various point numbers. Intuited by [43], we propose a novel way of contour representations based on the spherical coordinate system as shown in Fig.2. Under that circumstance, structural contour descriptors are efficiently calculated by recording the radial distance. Additionally, spherical centroids are proposed to enhance the precision of contour representations, serving as the spherical center. Then, vertebrae show similar 3D shapes according to anatomical structures, lumbar and thoracic vertebrae in particular. Hence, a low-dimension contour space can be extracted to represent all vertebrae data. Via a linear combination between low-rank basic descriptors in the contour space, instance masks of vertebrae can be approximately restored. Further, vertebral labels provided by coarse predictions facilitate the consistent generation of semantic masks. Ultimately, the whole pipeline is incorporated into SLoRD to boost the framework’s generalization abilities on unseen datasets. And the explicit shape constraint and basic contour descriptors are exerted to promote the regression precision of contours. SLoRD serves as a plug-and-play framework, which can repair the segmentation inconsistency from coarse predictions.
Our contributions are summarized as follows: (1) To achieve intra-vertebrae voxel consistency and contour precision simultaneously, we first propose the contour-based generation network named SLoRD, to address the multi-class vertebrae segmentation from CT images with different FOVs. (2) For a structural and efficient contour representation, the spherical coordinate system is adopted to describe the complete contour point cloud. Besides, spherical centroids are devised as the spherical center to boost the precision of contour representations, based on the anatomical characteristics of vertebrae. (3) Profiting from low-rank basic descriptors in the contour space, regressed contour points will be facilitated close to vertebral surfaces with shape and contour constraints. (4) Extensive experiments on VerSe 2019 and 2020 datasets demonstrate the superior performance of our methods compared with other single-stage and multi-stage approaches, with better segmentation consistency and boundary precision. Moreover, the SLoRD framework serves as a plug-and-play framework to repair intra-vertebrae segmentation inconsistency significantly.
@@
2 Related work
While the shape of individual vertebrae varies considerably along the spine, adjacent vertebrae still show fairly similar appearances, making them difficult to distinguish [29, 46]. Consequently, the spine presents unique challenges for quantitative image analysis compared to other anatomical structures. Additionally, factors such as pathological appearances, fractures, and metal implant artifacts [45] will introduce further challenges. Earlier studies [36, 16, 1, 2], cannot apply to the segmentation task of CT scans displaying arbitrary FOVs with various label distributions [42, 27]. To resolve these challenges, researchers resort to leveraging anatomical prior knowledge. Relevant research could be roughly categorized into two groups: single-stage and multi-stage frameworks for multi-class vertebrae segmentation.
For single-stage pipelines, Zheng et al. employed nnUNet [15] to perform multi-class vertebrae segmentation. Chang et al. [6] adopted stacked graph convolutional networks by featuring a pre-defined adjacency matrix to encode structural information. However, such approaches fail to model global dependency due to the intrinsic property of convolution [12]. Thus, Verteformer [45] was devised by fusing local details and global contexts with two parallel Transformers. And an edge detection block was proposed to refine intra-vertebrae boundaries. Nevertheless, it is still plagued by similar shapes between adjacent vertebrae and the patch-based inference paradigm, leading to intra-vertebrae segmentation inconsistency. Lessmann et al. [22] used fully convolutional networks by iteratively analyzing image patches for automatic vertebrae segmentation and identification. However, this method relies on the perfect prediction for the first label and hardly recovers from failures in the top vertebra.
In contrast, multi-stage approaches reveal more promising segmentation performance owing to the additional guidance, such as the heatmap and box-based localization of vertebrae. Sekuboyina et al. [34] proposed an interactive spine processing framework, consisting of spine detection by a light-weight network, vertebra labeling based on Btrfly Net [35], and vertebral segmentation by 3D UNet [11]. Payer et al. [31] implemented a three-step approach to first localize the spine, then simultaneously locate and identify each vertebra, and finally segment each vertebra. Besides, Tao et al. [38] introduced a Transformer to handle the labeling task, then employed UNet with the original CT and an auxiliary heatmap as inputs to segment all vertebrae. Chen et al. [7] devised a deep reasoning module leveraging anatomical prior knowledge to achieve anatomically correct results regarding the sequence of vertebrae. However, the multi-stage pipelines are prone to accumulating errors. An imprecise detected patch with inaccurate center regression will degrade the segmentation performance, particularly for pathological vertebrae with abnormal shapes or metal implants. As a result, the poor contour delineation will enforce inconsistency inside vertebrae.
@@ 3 Methodology
To simultaneously impose the voxel consistency and contour precision regularizations, we propose SLoRD by labeling sequential and complete binary masks to achieve multi-class vetebrae segmentation. Specifically, a spherical-based representation mode and spherical centroid are devised for structural and efficient contour representations as described in Section 3.1 and 3.2. Section 3.3 provides the details of SLoRD, which generates precise contour points based on explicit basic descriptors and contour regularization. The pipeline of the inference stage is illustrated in Section 3.4.
3.1 Spherical-based Contour Descriptors
Although the number of contour points varies across different vertebrae, a fixed number of sampling points is required for contour representations. However, it is quite tricky to sample contour points uniformly while maintaining the potential to restore the original contour. Thus, we incorporate the idea from PolarMask [43, 30], in which each contour point is encoded with the polar coordinate.
Figure 4: The two-stage framework. (a) Coarse predictions by arbitrary segmentation networks. (b) The generation of Gaussian positional prompts to boost the center localization precision. (c) Based on the combination of basic contour descriptors, SLoRD aims to regress contour points close to vertebral boundaries via explicit contour and shape regularizations.
Given the difficulty in extending the polar system to contour representations of 3D objects like vertebrae, we formulate contours with the spherical coordinate system as shown in Fig.2. Conditioning the inner centroid as the spherical center, we describe the boundary using spherical coordinates $(\rho_{{}_{\mathrm{ij}}},\theta_{i},\phi_{j}),i\mbox{=}1,...,I;j\mbox{=}1,...,J$. Here $\rho$ represents the length of the radial vector from the spherical center to the contour point, $\phi$ is the azimuth angle between the radial vector and $z^{+}$, $\theta$ means the polar angle between $x^{+}$ and the projected radial vector in the XOY plane. Angular coordinates $\theta_{i}$ and $\phi_{j}$ are sampled uniformly with a sampling interval $s$, with $I$ equal to $360/s$, $J$ equal to $(180/s\mbox{+}1)$. Thus, given the sampling angle grid, only radial coordinates are recorded to represent the contour:
@@ 3.1 Spherical-based Contour Descriptors
Here $\bm{\rho}$ refers to the contour descriptor, with the dimension $N$ equal to $I\times J$. Considering vertebral anatomies with the non-convex shape property, the spherical system chooses the maximum distance at the same angle to strike a balance between precision and computational cost. Since the shapes of most vertebrae are well structured and resembling to each other, low-rank contour descriptors can be exploited in a data-driven manner via Singular Value Decomposition (SVD) [30]. Detailedly, a non-convex contour matrix $\mathcal{M}=[\bm{\rho}_{1},\cdot\cdot\cdot,\bm{\rho}_{L}]$ is constructed from L vertebrae in the training data. Then we perform SVD on matrix $\mathcal{M}$:
@@
where $\bm{U}^{N\times N}=[\bm{u}_{1},\cdot\cdot\cdot,\bm{u}_{N}]$ and $\bm{V}^{L\times L}=[\bm{v}_{1},\cdot\cdot\cdot,\bm{v}_{L}]$ are orthogonal matrixes and $\Sigma^{N\times L}$ is a rectangular diagonal matrix, consisting of singular values $[\sigma_{1},\cdot\cdot\cdot,\sigma_{r}]$ with r equal to $\min(N,L)$. The best rank-k approximation of $\mathcal{M}$ can be calculated as follows:
$\mathcal{M}(k)$ $=\sigma_{1}\bm{u}_{1}\bm{v}_{1}^{T}+\cdot\cdot\cdot+\sigma_{k}\bm{u}_{k}\bm{v}_{k}^{T}$ (3)
$=[\bm{u}_{1},\cdot\cdot\cdot,\bm{u}_{k}]\,\mathcal{C}=\bm{U}(k)\,\mathcal{C}\approx\mathcal{M}$
$\mathcal{C}=[\sigma_{1}\bm{v}_{1},\cdot\cdot\cdot,\sigma_{k}\bm{v}_{k}]^{T}=\bm{U}(k)^{T}\mathcal{M}$
Here $\bm{U}(k)^{N\times k}$ refers to k basic contour descriptors, which can span into the contour space and be acquired from the training data in advance. $\mathcal{C}^{k\times L}$ means combinational coefficients, calculated by eigen descriptors and matrix $\mathcal{M}$. As long as $\mathcal{C}$ is correct, we can restore the original contours of vertebrae approximately as illustrated by Fig.3.
Another relevant research DeepSSM [4], as a shape model based on principal component analysis (PCA), aims to learn the functional mapping from images to low-dimensional latent codes. Then the model achieves discrete contour representations by applying the linear projection on these codes with the PCA basis and mean shape. In comparison, SVD bears the best low-rank approximation properties, and extracted basic descriptors are more robust to fit various anatomical shapes than PCA [40]. Also, the sampling strategy in [4] for the contour point cloud struggles to perform well for pathological vertebral structures with more diverse shapes and topology.
@@ 3.2 Spherical Centroid
Although the encoding scheme for vertebral contours is efficient, it is hard to determine the spherical center, which will significantly affect the restoration performance of vertebral shapes. Since the azimuth angle $\phi$ and polar angle $\theta$ are both sampled at equal intervals, a spherical center with a biased location will result in fewer sampling contour points for the boundary region far from it. Consequently, that region suffers from a contour under-representation. Hence, it has been a crucial topic on how to designate the spherical center.
@@
$&\mathcal{B}=\mathcal{F}(M)\setminus M\\ &C_{s}=\mathop{\arg\min}\limits_{c\in M}(\tfrac{1}{|\mathcal{B}|}\displaystyle\sum_{b\in\mathcal{B}}d\,(c,b)+\lambda\cdot d\,(c,\delta)),\\ &\mathrm{s.t.}\ \delta=\mathop{\arg\max}\limits_{j\in\mathcal{B}}j_{y}\end{aligned}$ (4)
@@
where $\mathcal{F}(\cdot)$ represents the morphological dilation operation, $M$ refers to the segmentation mask of each vertebral volume, the boundary map $\mathcal{B}$ is the difference set between the dilated mask $\mathcal{F}(M)$ and the mask itself $M$. Besides, $d(\cdot)$ means the Euclidean distance. The first item in Eq (4) refers to the average distance from each boundary point $b$ to the center $c\in\mathcal{P}$. Concerning the non-convex shape prior of the spinous process, we attempt to push the spherical centroid closer to it. That will enlarge the value range of $\theta$ and $\phi$, then enhance the shape representations of the spinous process. That is why the second regularization term is incorporated, in which $\delta$ is the boundary point $j\in\mathcal{B}$ with the largest y-dimensional index $j_{y}$. $\lambda$ is a scaling parameter, empirically set as $0.005$.
3.3 Contour Generation with Explicit Contour Priors